OpenStudy (anonymous):

Find a if b = 60 and c = 61. Right triangle with sides a, b, and c. Side c is opposite the right angle. A. a = 121 B. a = 1 C. a = 11 D. a = the square root of7321

jimthompson5910 (jim_thompson5910):

We have a right triangle. So we can use the Pythagorean Theorem to solve for 'a'

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 a^2 + 60^2 = 61^2 a^2 + 3600 = 3721 ... ... ... a = ???

OpenStudy (anonymous):

So it would be B?

jimthompson5910 (jim_thompson5910):

a^2 + 3600 = 3721 a^2 + 3600 -3600 = 3721 - 3600 a^2 = 121 a = ???

OpenStudy (anonymous):

A? .-.

jimthompson5910 (jim_thompson5910):

that's a^2, not 'a'

jimthompson5910 (jim_thompson5910):

what undoes squaring?

OpenStudy (anonymous):

Would you subtract?

jimthompson5910 (jim_thompson5910):

what operation is the opposite of squaring?

OpenStudy (anonymous):

Square root?

jimthompson5910 (jim_thompson5910):

good, so take the square root of both sides

OpenStudy (anonymous):

It would be C?

jimthompson5910 (jim_thompson5910):

yes because \(\Large \sqrt{121} = 11\)

OpenStudy (anonymous):

Okay thanks so much. Do you mind helping with several more?

jimthompson5910 (jim_thompson5910):

I can do a few more, sure

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

An airplane lands at an airport 60 miles east and 25 miles north of where it took off. How far apart are the two airports? image of a triangle with a box in the lower right corner; the horizontal side forming that corner is labeled 60 miles and the vertical side forming that corner is labeled 25 miles A. 65 miles B. 54.54 miles C. 13.04 miles D. 9.22 miles

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 60^2 + 25^2 = x^2 solve for x

OpenStudy (anonymous):

So.. 120+50=c2 ?

jimthompson5910 (jim_thompson5910):

60^2 means 60 squared NOT 60 times 2

OpenStudy (anonymous):

Oh goodness I knew that. My apologies. So 3600+625=c2

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

now add to get _____

OpenStudy (anonymous):

4225?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so \[\Large x^2 = 4225\] how do we isolate x?

OpenStudy (anonymous):

I'm not to sure.

jimthompson5910 (jim_thompson5910):

we do the same thing we did back when we solved for x in \(\Large x^2 = 121\)

OpenStudy (anonymous):

So it would be 65?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

With your help of course. Thank you for taking the time to explain it.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Find c if a = 5 and b = 12. Right triangle with sides a, b, and c. Side c is opposite the right angle. A. c = 17 B. c = the square root of17 C. c = 13 D. c = the square root of13

OpenStudy (anonymous):

So this would be 25+144=169=13

jimthompson5910 (jim_thompson5910):

you are correct, but the notation and the way you wrote it is a bit off you should have \(\Large \sqrt{169} = 13\) but you have the right idea

OpenStudy (anonymous):

Yeah I just didn't know how to put that.

jimthompson5910 (jim_thompson5910):

oh you can type sqrt to mean "square root" so sqrt(169) is the square root of 169

OpenStudy (anonymous):

Therefore this would be B? Find c if a = 3 and b = 5. Right triangle with sides a, b, and c. Side c is opposite the right angle. A. c = 4 B. c = the square root of34 C. c = 8 D. c = the square root of15

OpenStudy (anonymous):

Thank you.

jimthompson5910 (jim_thompson5910):

B) \(c = \sqrt{34}\) is correct

OpenStudy (anonymous):

YAAAAY!

jimthompson5910 (jim_thompson5910):

sounds like you got the hang of this, that's good

OpenStudy (anonymous):

Thanks so much for helping again!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

Actually, I have 1 more?

OpenStudy (anonymous):

Which of the following equations does not represent the Pythagorean Theorem? image of a triangle whose sides are labels a, b, and c; the corner of sides a and b have a box A. a2 + b2 = c2 B. a2 - b2 = c2 C. a2 = c2 – b2 D. c2 – a2 = b2

jimthompson5910 (jim_thompson5910):

which one do you think it is?

jimthompson5910 (jim_thompson5910):

I'll start you off. The Pythagorean Theorem is \[\Large a^2 + b^2 = c^2\]

OpenStudy (anonymous):

I think B

jimthompson5910 (jim_thompson5910):

You can isolate \(\Large a^2\) to get this \[\Large b^2 = c^2-a^2\] then flip the equation to get \[\Large c^2-a^2 = b^2\]

jimthompson5910 (jim_thompson5910):

B is correct

OpenStudy (anonymous):

Find c if a = 3 and b = 4. Right triangle with sides a, b, and c. Side c is opposite the right angle. A. c = 7 B. c = 5 C. c = 7/2 D. c = the square root of7

OpenStudy (anonymous):

Sorry I'm confused again.

jimthompson5910 (jim_thompson5910):

\[\Large a^2 + b^2 = c^2\] \[\Large 3^2 + 4^2 = c^2\] solve for c

OpenStudy (anonymous):

sqrt(21)

OpenStudy (anonymous):

OMG fail.

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

I added wrong-.-

jimthompson5910 (jim_thompson5910):

B is correct

OpenStudy (anonymous):

Thanks.

jimthompson5910 (jim_thompson5910):

np